OpenStudy (anonymous):

find the volume of a pyramid with an equilateral triangle as a base that has a side length of 12 in. The height of the pyramid is 10 in

OpenStudy (anonymous):

Surface pyramid with an equilateral triangle as a base = 3 * surface triangle + surface equilateral triangle Surface triangle = 0.5bh b=12 in h=0.5*sqrt(3)*12 = 6sqrt(3) in Surface triangle = 0.5 * 12 * 6*sqrt(3)=36sqrt(3) in² Surface equilateral triangle = 0.5 * sqrt(3) * 12=6sqrt(3) in² Surface pyramid = 3 * 36sqrt(3) + 6sqrt(3) = 108sqrt(3) +6 sqrt(3)=114 sqrt(3) in²